# Deep Neural Networks: Multi-Classification and Universal Approximation

### v1

@@ 6. Further comments and open problems
According to Step 2.3 in the proof of Theorem 1.2, the map $\phi^{\mathcal{L}}=\phi^{L}\circ\phi^{2N_{E}}$ drives the hyperrectangles defined in $\mathcal{H}$ into their respectively labels. Therefore, to estimate the norm of $\phi^{\mathcal{L}}$, we divide the proof into two parts.
@@
1) Precondition of the data: In this phase, $b_{1}$ is chosen big enough such that the $\bm{\sigma}$ acts as the identity function. Considering as an input data point the output of the map $\phi^{2N_{E}}$, then it is enough to take $b$ bigger than $C(h+\delta/2^{2N_{E}})$. This implies:
@@
2) Compression process: After the Precondition of the data, all the datasets have been projected to the real line, and the distance between points does not exceed $C\left(\delta/2^{2N_{E}}\right)$. We place the hyperplanes in the “Compression process” depending on the data location. Therefore, we deduce that
@@
On the other hand, given $\hat{L}>0$ and a family of parameters $\mathcal{W}^{\hat{L}}=\{W^{i}\}_{i=1}^{\hat{L}}$ and $\mathcal{B}^{\hat{L}}=\{b^{i}\}_{i=1}^{\hat{L}}$, the norm of $\phi^{\hat{L}}:=\phi^{\hat{L}}(\mathcal{W}^{\hat{L}},\mathcal{B}^{\hat{L}},\cdot)$ the input-output map of (1.1) can be bounded by
$\|\phi^{\hat{L}}\|_{L^{\infty}(\mathcal{C};\mathbb{R}_{+})}\leq\underset{x\in\mathcal{C}}{\operatorname{supess}}\left\|\prod_{j=1}^{\hat{L}}W^{j}x+\sum_{i=1}^{\hat{L}-1}\left(\prod_{j=i}^{\hat{L}-1}W^{j+1}\right)b^{i}+b^{\hat{L}}\right\|_{\infty}$ (B.7)
@@
$\|\phi^{\mathcal{L}}\|_{L^{\infty}(\mathcal{C};\mathbb{R}_{+})}=\|\phi^{L}\circ\phi^{2N_{E}}\|\leq C\,\underset{x\in\mathcal{C}}{\operatorname{supess}}\|\phi^{2N_{E}}(x)\|\\ +2N_{h}C\left(\frac{\delta}{2^{2N_{E}}}\right)+(2M_{h}+1)C\left(\frac{\delta}{2^{2N_{E}}}\right)+2C\left(h+\frac{\delta}{2^{N_{E}}}\right).$
@@
Denote by $l_{\mathcal{C}}$ the largest edge of $\mathcal{C}$. Then, applying (5.28) in (B.8) there exists a positive constant $C_{1}$, independents of $h$ and $\delta$, such that
$\|\phi^{\mathcal{L}}\|_{L^{\infty}(\mathcal{C};\mathbb{R}_{+})}$ $\leq C_{1}\left(1+\delta\frac{2^{\frac{-C_{2}}{h+\delta}}}{(h+\delta)^{d}}+\delta 2^{\frac{-C_{2}}{h+\delta}}+h\right),$
where $C_{2}=2l_{\mathcal{C}}$. Applying the inequality $|a+c|^{p}\leq 2^{p}(|a|^{p}+|c|^{p})$ three times, we obtain:
$\|\phi^{\mathcal{L}}\|_{L^{\infty}(\mathcal{C};\mathbb{R}_{+})}^{p}\leq C_{1}^{p}2^{3p}\left(1+\delta^{p}\frac{2^{\frac{-pC_{2}}{h+\delta}}}{(h+\delta)^{dp}}+\delta^{p}2^{\frac{-C_{2}p}{h+\delta}}+h^{p}\right).$ (B.9)
Now, since $e^{y}\geq\sum_{k=0}^{dp+1}y^{k}/k!$ using the change of variable $y=log(2)C_{2}/(h+\delta)$ we deduce the inequality
$\frac{2^{\frac{-pC_{2}}{h+\delta}}}{{(h+\delta)^{dp}}}$ $\leq\frac{1}{(h+\delta)^{dp}}\left(\sum_{k=0}^{dp+1}\frac{(C_{2}\log(2))^{k}(h+\delta)^{-k}}{k!}\right)^{-1}$
$=\frac{1}{(h+\delta)^{dp}}\left(\frac{1}{(h+\delta)^{dp+1}}\sum_{k=0}^{dp+1}\frac{(C_{2}\log(2))^{k}(h+\delta)^{(dp+1)-k}}{k!}\right)^{-1}$
$=(h+\delta)\left(\frac{(C_{2}\log(2))^{dp+1}}{(dp+1)!}+\sum_{k=0}^{dp}\frac{(C_{2}\log(2))^{k}(h+\delta)^{(dp+1)-k}}{k!}\right)^{-1}.$
Therefore, for $h<\frac{l_{\mathcal{C}}\log(2)}{(dp+1)}$ we have that
$\frac{2^{\frac{-pC_{2}}{h+\delta}}}{{(h+\delta)^{dp}}}\leq\frac{(dp+1)!}{(2l_{\mathcal{C}}\log(2))^{dp+1}}(h+\delta),$
Similarly, using the inequality $e^{y}\geq 1+y$ and the change of variable $y=log(2)C_{2}/(h+\delta)$, we deduce that for $h<l_{\mathcal{C}}\log(2)$ we have that $2^{\frac{-pC_{2}}{h+\delta}}\leq(l_{\mathcal{C}}\log(2))^{-1}(h+\delta)$. Consequently, from (B.9) the obtain the inequality
$\|\phi^{\mathcal{L}}\|_{L^{\infty}(\mathcal{C};\mathbb{R}_{+})}^{p}\leq C\left(1+\delta^{p}(h+\delta)+h^{p}\right),$
with $C>0$ a constant depending on $d$, $m_{d}(\mathcal{C})$, $p$ and $\|f\|_{L^{p}(\Omega;\mathbb{R}_{+})}$. This concludes the proof. ∎
@@ Acknowledgments
The authors wish to express their gratitude to D. Ruiz-Balet for his insightful discussions and to A.Álvarez-López and T. Crin-Barat, for taking the time to critically review our manuscript.
M. Hernández has been funded by the Transregio 154 Project, Mathematical Modelling, Simulation, and Optimization Using the Example of Gas Networks of the DFG, project C07, and the fellowship ”ANID-DAAD bilateral agreement”. E. Zuazua has been funded by the Alexander von Humboldt-Professorship program, the ModConFlex Marie Curie Action, HORIZON-MSCA-2021-DN-01, the COST Action MAT-DYN-NET, the Transregio 154 Project Mathematical Modelling, Simulation and Optimization Using the Example of Gas Networks of the DFG, AFOSR 24IOE027 project, grants PID2020-112617GB-C22 and TED2021-131390B-I00 of MINECO (Spain), and grant PID2023-146872OB-I00 of MICIU (Spain). Madrid Government - UAM Agreement for the Excellence of the University Research Staff in the context of the V PRICIT (Regional Programme of Research and Technological Innovation). Both authors have been partially supported by the DAAD/CAPES Programs for Project-Related Personal, grant 57703041 ’Control and numerical analysis of complex system’.

### v2

@@ 7. Further comments and open problems
Let us begin by observing that Corollary 1.1 cannot be directly applied since now we have an infinite number of data points. However, we can provide a similar estimation by carefully analyzing the parameters used in the proof of Theorem 1.4. According to Step 2.3 in the proof of Theorem 1.4, the map $\phi^{\mathcal{L}}=\phi^{L}\circ\phi^{2N_{E}}$ drives the hyperrectangles defined in $\mathcal{H}$ into their respectively labels. Therefore, to estimate the norm of $\phi^{\mathcal{L}}$, we divide the proof into two parts.
@@
1) Precondition of the data: In this phase, $b_{1}$ is chosen large enough such that $\bm{\sigma}$ acts as the identity function. Considering as an input data point the output of the map $\phi^{2N_{E}}$, then it is enough to take $b$ larger than $C(h+\delta/2^{2N_{E}})$. This implies:
@@
2) Compression process: After data preconditioning, all the datasets have been projected to the real line, and the distance between points does not exceed $C\left(\delta/2^{2N_{E}}\right)$. We place the hyperplanes in the “Compression process” depending on the data location. Therefore, we deduce that
@@
On the other hand, given $\hat{L}>0$ and a family of parameters $\mathcal{W}^{\hat{L}}=\{W^{i}\}_{i=1}^{\hat{L}}$ and $\mathcal{B}^{\hat{L}}=\{b^{i}\}_{i=1}^{\hat{L}}$, the norm of $\phi^{\hat{L}}:=\phi^{\hat{L}}(\mathcal{W}^{\hat{L}},\mathcal{B}^{\hat{L}},\cdot)$, the input-output map of (1.2), can be bounded by
$\|\phi^{\hat{L}}\|_{L^{\infty}(\mathcal{C};\mathbb{R}_{+})}\leq\underset{x\in\mathcal{C}}{\operatorname*{ess\,sup}}\left\|\prod_{j=1}^{\hat{L}}W^{j}x+\sum_{i=1}^{\hat{L}-1}\left(\prod_{j=i}^{\hat{L}-1}W^{j+1}\right)b^{i}+b^{\hat{L}}\right\|_{\infty}$ (B.7)
@@
$\|\phi^{\mathcal{L}}\|_{L^{\infty}(\mathcal{C};\mathbb{R}_{+})}=\|\phi^{L}\circ\phi^{2N_{E}}\|\leq C\,\underset{x\in\mathcal{C}}{\operatorname*{ess\,sup}}\|\phi^{2N_{E}}(x)\|\\ +2N_{h}C\left(\frac{\delta}{2^{2N_{E}}}\right)+(2M_{h}+1)C\left(\frac{\delta}{2^{2N_{E}}}\right)+2C\left(h+\frac{\delta}{2^{N_{E}}}\right).$
@@
Denote by $l_{\mathcal{C}}$ the largest edge of $\mathcal{C}$. Then, applying (5.28) in (B.8) there exists a positive constant $C_{1}$, independent of $h$ and $\delta$, such that
$\|\phi^{\mathcal{L}}\|_{L^{\infty}(\mathcal{C};\mathbb{R}_{+})}$ $\leq C_{1}\left(1+\delta\frac{2^{\frac{-C_{2}}{h+\delta}}}{(h+\delta)^{d}}+\delta 2^{\frac{-C_{2}}{h+\delta}}+h\right),$ (B.9)
where $C_{2}=2l_{\mathcal{C}}$. Now, since $e^{y}\geq\sum_{k=0}^{d+1}y^{k}/k!$ using the change of variable $y=log(2)C_{2}/(h+\delta)$ we deduce that
$\frac{2^{\frac{-C_{2}}{h+\delta}}}{{(h+\delta)^{d}}}$ $\leq\frac{1}{(h+\delta)^{d}}\left(\sum_{k=0}^{d+1}\frac{(C_{2}\log(2))^{k}(h+\delta)^{-k}}{k!}\right)^{-1}$
$=\frac{1}{(h+\delta)^{d}}\left(\frac{1}{(h+\delta)^{d+1}}\sum_{k=0}^{d+1}\frac{(C_{2}\log(2))^{k}(h+\delta)^{(d+1)-k}}{k!}\right)^{-1}$
$=(h+\delta)\left(\frac{(C_{2}\log(2))^{d+1}}{(d+1)!}+\sum_{k=0}^{d}\frac{(C_{2}\log(2))^{k}(h+\delta)^{(d+1)-k}}{k!}\right)^{-1}.$
Therefore, for $h<\frac{l_{\mathcal{C}}\log(2)}{(d+1)}$ we have that
$\frac{2^{\frac{-C_{2}}{h+\delta}}}{{(h+\delta)^{d}}}\leq\frac{(dp+1)!}{(2l_{\mathcal{C}}\log(2))^{d+1}}(h+\delta),$
Similarly, using the inequality $e^{y}\geq 1+y$ and the change of variable $y=log(2)C_{2}/(h+\delta)$, we deduce that for $h<l_{\mathcal{C}}\log(2)$ we have that $2^{\frac{-C_{2}}{h+\delta}}\leq(l_{\mathcal{C}}\log(2))^{-1}(h+\delta)$. Consequently, from (B.9) we obtain the inequality
$\|\phi^{\mathcal{L}}\|_{L^{\infty}(\mathcal{C};\mathbb{R}_{+})}\leq C\left(1+\delta(h+\delta)+h\right),$
with $C>0$ a constant depending on $d$, $m_{d}(\mathcal{C})$, and $\|f\|_{L^{p}(\Omega;\mathbb{R}_{+})}$. This concludes the proof. ∎
@@ Acknowledgments
The authors wish to express their gratitude to D. Ruiz-Balet and A. Alcalde for their insightful discussions and to A.Álvarez-López and T. Crin-Barat, for taking the time to critically review our manuscript.
M. Hernández has been funded by the Transregio 154 Project, Mathematical Modelling, Simulation, and Optimization Using the Example of Gas Networks of the DFG, project C07, and the fellowship ”ANID-DAAD bilateral agreement”.
E. Zuazua was funded by the European Research Council (ERC) under the European Union’s Horizon 2030 research and innovation programme (grant agreement NO: 101096251-CoDeFeL), the Alexander von Humboldt-Professorship program, the ModConFlex Marie Curie Action, HORIZON-MSCA-2021-dN-01, the COST Action MAT-DYNNET, the Transregio 154 Project of the DFG, AFOSR Proposal 24IOE027 and grants PID2020-112617GB-C22/AEI/10.13039/501100011033 and TED2021131390B-I00/ AEI/10.13039/501100011033 of MINECO (Spain), and Madrid GovernmentUAM Agreement for the Excellence of the University Research Staff in the context of the V PRICIT (Regional Programme of Research and Technological Innovation).
Both authors have been partially supported by the DAAD/CAPES Programs for Project-Related Personal, grant 57703041 ’Control and numerical analysis of complex system’.